# AC3D: Analyzing and Improving 3D Camera Control in Video Diffusion Transformers

### v2

@@ Appendix D Motion analysis details
Appendix E Linear probing details
@@
Appendix F Dataset construction details
As describe in Section 3.5, we construct a dataset of videos with scene motion but recorded from stationary cameras. One might attempt to build such a dataset in an automated way by estimating the optical flow and checking whether there is non-zero motion in the middle region of the frame, and no motion on the borders. However, this would not work well for a myriad of corner cases, which is why we opted for manual construction. For this, we annotated 110K internal videos for the presence of camera and the presence of scene motion. Each video has a duration of $5-30$ seconds and covers various diverse categories of scenes: humans, animals, landscapes, food and other types. Out of these videos, 20K videos turned out to be the necessary ones: with scene motion, but completely static scenes. We proceeded to use them as our training data without further processing. As being discussed in Appendix C, we augmented their camera information artifically by using random extrinsics from RealEstate10K [197].
Appendix G Scaling bias
@@
Appendix H Additional Related Work

### v4

@@ Appendix E Motion analysis details
Appendix F Linear probing details
@@
Appendix G Dataset construction details
As describe in Section 3.5, we construct a dataset of videos with scene motion but recorded from stationary cameras. One might attempt to build such a dataset in an automated way by estimating the optical flow and checking whether there is non-zero motion in the middle region of the frame, and no motion on the borders. However, this would not work well for a myriad of corner cases, which is why we opted for manual construction. For this, we annotated 110K internal videos for the presence of camera and the presence of scene motion. Each video has a duration of $5-30$ seconds and covers various diverse categories of scenes: humans, animals, landscapes, food and other types. Out of these videos, 20K videos turned out to be the necessary ones: with scene motion, but completely static scenes. We proceeded to use them as our training data without further processing. As being discussed in Appendix D, we augmented their camera information artifically by using random extrinsics from RealEstate10K [201].
Appendix H Scaling bias
@@
Appendix I Additional Related Work
@@
Appendix J User Study
  • Motion Quality (MQ): Which video has larger and more natural motion.
  • Visual Quality (VQ): Which video has a higher overall visual quality.